# Assessing Inf luential Observations in Pain Prediction using fMRI Data

### v1

@@ 5.3 Analysis of Real Datasets
    Different combinations of influential observation detection and model selectors lead to drastically distinctive outcomes. Specifically, we can see that while no points are identified as influential by ClusMIP(SLASSO) and MIP, no less than 16 points are labelled influential by ClusMIP(LASSO), ClusMIP(ENET), ClusMIP(SCAD) and ClusMIP(MCP). This substantiates our previous explanation in Section 2 that the detection of influential observations in our context is intricately linked to the properties of variable selection procedures. To be more specific, different penalty functions ingrained in the model selectors reveal different sensitivities with respect to variable selection in the presence of potential influential observations. For this particular dataset, LASSO seems to be the most sensitive model selector, followed by SCAD and MCP. Furthermore, the matching performance of ClusMIP(SLASSO) and MIP aligns with our simulation results in Section 4.4, where $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$ and $\text{Power}_{\text{MIP}}$ are similar under certain scenarios.
    Table 1: Assessment Metrics for the Pain Prediction Dataset (Lindquist et al., 2017)
    Detection(Selectors)
    $\widehat{\text{I}}^{\text{infl}}_{\text{Detection(Selector)}}$
    $\big{|}\widehat{\text{I}}^{\text{infl}}_{\text{Detection (\text{Selector}})}\big{\rvert}$
    $\widehat{\text{T}}_{\text{Detection(Selector)}}$
    ClusMIP(LASSO)
    3, 7, 13, 14, 15, 19, 22, 25, 26, 43, 44, 49, 50, 52, 55, 56, 57, 58, 62, 67, 75, 79, 80, 81, 103, 115, 122, 123, 124, 127, 145, 147, 157, 158, 159, 164, 165, 169, 170, 172, 175, 176, 194
    42
    3214.03
    ClusMIP(SLASSO)
    N.A.
    0
    2103.24
    ClusMIP(ENET)
    31, 49, 50, 52, 55, 58, 76, 79, 115, 122, 124, 158, 169, 170, 172, 194
    16
    45944.46
    ClusMIP(SCAD)
    31, 43, 49, 52, 55, 57, 62, 76, 80, 86, 103, 121, 133, 134, 145, 146, 157, 159, 164, 165, 173, 177, 181, 182, 193
    25
    6474.15
    ClusMIP(MCP)
    13, 19, 26, 49, 52, 55, 57, 58, 73, 76, 86, 103, 121, 124, 134, 145, 147, 151, 163, 164, 165, 182
    22
    5059.49
    MIP
    N.A.
    0
    5.27
    DF(LASSO)
    2, 57, 84, 119, 153, 159, 168, 174, 198
    9
    64.83
  2. 2.
@@
    For ClusMIP(LASSO) and ClusMIP(ENET), we can see that $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(ENET)}}$ is almost a proper subset of $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(LASSO)}}$, except for observations 31 and 76. Indeed, $\lvert\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(LASSO)}}\rvert$ is the largest among all procedures considered. One possible explanation is that for this dataset, predictors which are multivariate brain activity patterns represented by voxels are expected to be highly correlated. As such, certain significant predictors may be de-selected by the LASSO. On the other hand, the unification of both $\ell_{1}$ and $\ell_{2}$ penalties embedded in the ENET facilitates the improved performance in variable selection. Thus, the submodel selected by LASSO may be more parsimonious than the one selected by ENET. This necessarily implies that in the finite sample, the submodel selected by ENET may demonstrate better goodness-of-fit than the one chosen by LASSO. Therefore, in the presence of influential observations, LASSO may have a higher tendency to be impacted by perturbation in variable selection than the ENET.
  3. 3.
    $\widehat{\text{I}}^{\text{infl}}_{\text{DF}(\text{LASSO})}$ hardly overlaps with those identified by other methods, except for the observations 57 and 169. This is mainly due to the embedded leave-one-out design, implying its inability to accurately capture multiple influential points.
@@
    Figure 5: (A): predictive power of the full and reduced datasets; (B): pain scores as a function of $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(MCP)}}$; (C): the first two dimensions from multi-dimensional scaling.
  4. 4.
@@
    We have $\widehat{\text{I}}^{\text{infl}}_{\text{Common}}\coloneqq\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{LASSO})}\ \cap\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{ENET})}\ \cap\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{SCAD})}\ \cap\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{MCP})}=\{49,52,55\}$. As such, observations 49, 52 and 55 may serve as observations of interest that could undergo further studies. The first two observations correspond to two different temperatures measured on the same subject, indicating this may be a subject that needs to be further investigated. In addition, the first and third observation correspond to the lowest temperature, possibly indicating subjects that have a differing pain threshold than other subjects. Indeed, graphical illustration of pain ratings as a function of $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(Selector)}}$ (Figure 5(B)) indicates that the majority correspond to low pain ratings under each designated temperature, a pattern which holds over all the model selectors. In fact, the low pain score primarily contributes to the “outlyingness” of the identified influential points, since no anomalies in the feature space are detected as seen in the multi-dimensional scaling plot (Figure 5(C)). This may be explained by the fact that this temperature is generally considered warm, but not painful, thus differentiating it from other temperature levels.
  5. 5.
@@
    Different model selection outcomes are observed when a submodel is selected from the full dataset and the one after removing $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(Selector)}}$. In Figure 6, we highlight the brain regions selected via MCP with color intensities represented by the corresponding coefficient estimates obtained from the full and reduced datasets. Panels A and B demonstrate such discrepancy in selected regions, while Panels C, D and E pinpoint common and uniquely selected regions from the two datasets. Six regions are common to both models, including primary visual, frontoparietal, dorsal attention, and default mode areas, as well as the amygdala. Regions unique to the full model include the cerebellum, while regions unique to the reduced model include ventral attention and additional frontoparietal areas.
@@
  6. 6.
    In Figure 5 (A), we observe improved sample correlation between the real and predicted pain scores after removing $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(Selector)}}$ and $\widehat{\text{I}}^{\text{infl}}_{\text{Common}}$ for all the model selectors. For example, such correlation increases from 0.53 to 0.80 after removing $\widehat{\text{I}}^{\text{infl}}_{\text{MCP}}$, illustrating the advantage for identifying and removing influential points.
6 Conclusion
In this manuscript, we highlight the important issue of identifying observations that could unduly affect the choice of a stochastically selected submodel, which is largely ignored in the existing influential diagnosis literature. Indeed, inappropriate handling of such observations unavoidably leads to an inaccurate selected submodel (Figure 1 in Section 1 of the Supplementary Materials), giving rise to detrimental effects on the validity and reproducibility in downstream statistical inference. To capture them, especially under the high-dimensional setting, we have proposed the so-called ClusMIP detection procedure, and evaluated it alongside two existing proposals, the DF(LASSO) and MIP. From the simulation results in Section 4, we see that the DF(LASSO) is unable to detect multiple influential points due to its leave-one-out design. On the other hand, the MIP is a computationally efficient and effective diagnostic metric in dealing with contamination introduced only on the response, and on the response and predictors simultaneously. Yet, it under-performs when contamination is solely included in the predictors. In comparison, the performance of the ClusMIP varies according to data generating mechanisms and model selectors. In fact, the variation in the performance of ClusMIP as a function of model selector provides insights on the sensitivities of different model selectors, which is one essential objective of this manuscript. To be more precise, under the linear regression model and when predictors are independent, ClusMIP(SLASSO) offers relatively satisfactory solutions in detecting all the influential points, implying that both model selectors are sensitive to their presence. In contrast, other model selectors may not be sensitive as shown in their detection powers. Yet, one limitation of the ClusMIP procedure lies in its computational cost, which depends on the outcome of the selected high-dimensional clustering procedure.
In addition to the above conclusions, we now outline possible extensions of the current work. First, we recognize significant variation in the detection performances of the ClusMIP algorithm on different model selectors. Initial investigation reveals that in the second stage of the procedure, the GDF measure (5) associated with some selectors may not be sufficiently large to recognize true influential points. Yet, the underlying theoretical logic, which is intricately linked to the breakdown points for model selectors that are briefly discussed in Section 2, is unclear and currently undergoing further investigation. Second, the exact finite- and large-sample distributional properties of $\tau_{i}$ in (5) have not been studied, which may pave the way for alternative parametric and re-sampling approximation.
{funding}
Dongliang Zhang is supported in part by NIH grant R01MH129397 from the National Institute of Mental Health (NIMH). Masoud Asgharian is supported by the Natural Science and Engineering Research Council of Canada (NSERC RGPIN-2018-05618). Martin A. Lindquist is supported in part by NIH grant R01 EB026549 from the National Institute of Biomedical Imaging and Bioengineering (NIBIB), and R01MH129397 from NIMH.
[…]
The Supplementary Materials referenced in Sections 1 to 5 contains data generating details of the motivating example, proofs of Theorems 1 to 3 and Corollary 1, algorithmic details of the RGD, MIP and RGDF procedures, comparison of different clustering procedures, and additional simulation details and plots.

### v2

@@ 6.1 Detection of Influential Observations
From Table 1, we observe that different combinations of detection procedures and model selectors cause drastically distinctive detection outcomes. Specifically, while no points are labeled as influential by ClusMIP(SLASSO) and MIP, more than 16 observations are deemed influential by ClusMIP(LASSO), ClusMIP(ENET), ClusMIP(SCAD) and ClusMIP(MCP). This substantiates our previous explanation in Section 3 that the detection of influential observations on variable selection is intricately linked to the properties of variable selection procedures. To be more specific, different penalty functions of the model selectors reveal different sensitivities with respect to variable selection in the presence of influential observations. For this particular dataset, LASSO seems to be the most sensitive model selector, followed by SCAD, MCP and ENET.
Table 1: Assessment Metrics for the Pain Prediction Dataset [21]
Detection(Selectors) $\widehat{\text{I}}^{\text{infl}}_{\text{Detection(Selector)}}$ $\big{|}\widehat{\text{I}}^{\text{infl}}_{\text{Detection (\text{Selector}})}\big{\rvert}$
ClusMIP(LASSO) 3, 7, 13, 14, 15, 19, 22, 25, 26, 43, 44, 49, 50, 52, 55, 56, 57, 58, 62, 67, 75, 79, 80, 81, 103, 115, 122, 123, 124, 127, 145, 147, 157, 158, 159, 164, 165, 169, 170, 172, 175, 176, 194 42
ClusMIP(SLASSO) N.A. 0
ClusMIP(ENET) 31, 49, 50, 52, 55, 58, 76, 79, 115, 122, 124, 158, 169, 170, 172, 194 16
ClusMIP(SCAD) 31, 43, 49, 52, 55, 57, 62, 76, 80, 86, 103, 121, 133, 134, 145, 146, 157, 159, 164, 165, 173, 177, 181, 182, 193 25
ClusMIP(MCP) 13, 19, 26, 49, 52, 55, 57, 58, 73, 76, 86, 103, 121, 124, 134, 145, 147, 151, 163, 164, 165, 182 22
MIP N.A. 0
DF(LASSO) 2, 57, 84, 119, 153, 159, 168, 174, 198 9
@@
For ClusMIP(LASSO) and ClusMIP(ENET), we observe that $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(ENET)}}$ is almost a proper subset of $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(LASSO)}}$, except for observations 31 and 76. Indeed, $\lvert\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(LASSO)}}\rvert$ is the largest among all procedures considered. One possible explanation is that for this dataset, predictors which are multivariate brain activity patterns represented by voxels are expected to be highly correlated. Therefore, certain significant predictors may fail to be selected by LASSO. On the other hand, the use of both $\ell_{1}$ and $\ell_{2}$ penalties in ENET facilitates the improved performance in variable selection. Thus, the submodel selected by LASSO may be more parsimonious than the one selected by ENET. This necessarily implies that in the finite sample, the submodel selected by ENET may demonstrate better goodness-of-fit than the one chosen by LASSO. As such, in the presence of influential observations, LASSO may have a higher tendency to be impacted by perturbation in variable selection than ENET. Finally, the set $\widehat{\text{I}}^{\text{infl}}_{\text{DF}(\text{LASSO})}$ exhibits the lowest cardinality and hardly overlaps with those identified by other procedures, except for observation 57. In fact, this low cardinality is consistent with the simulation studies, where the detection power of DF(LASSO) is always among the lowest (Panel (A) of Figures 1 and 2).
@@
Figure 3: (A) and (B): respective distributions of $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{MCP})}$ and $\widehat{\text{I}}^{\text{infl}}_{\text{DF(LASSO)}}$. (C) and (D): the first two dimensions from multi-dimensional scaling respectively from the ClusMIP(MCP) and DF(LASSO).
@@
Furthermore, we observe that $\widehat{\text{I}}^{\text{infl}}_{\text{Common}}\coloneqq\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{LASSO})}\ \cap\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{ENET})}\ \cap\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{SCAD})}\ \cap\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{MCP})}=\{49,52,55\}$. Hence, observations 49, 52 and 55 are flagged for further investigation. The first two observations correspond to two different temperatures measured on the same participant, indicating this may be a participant that needs to be further investigated. In addition, the first and third observations correspond to the lowest temperature, possibly indicating participants that have a differing pain threshold than other participants. Indeed, a plot of the pain ratings as function of temperature with influential points identified by the ClusMIP(MCP) (Figure 3(A)) indicates that the majority of outliers correspond to low pain ratings at each temperature, a pattern which holds for ClusMIP across all model selectors. For illustration, only the plot for ClusMIP(MCP) is shown (Figure 3(A)) while the other plots in the ClusMIP family are presented in Section 12 of the Supplementary Materials Part I. In fact, the low pain score primarily contributes to the “outlyingness” of the identified influential points, as no anomalies in the feature space are detected as seen in the multi-dimensional scaling plot (Figure 3(C)). This may be explained by the fact that this temperature is generally considered warm, but not always painful, thus differentiating it from other temperature levels. Together these two panels imply that the low pain ratings may need to be described using a different model than that of high pains in fMRI-based thermal pain modeling. This assertion is further strengthened by research on pain processing mechanisms suggesting a discrepancy in brain activity between low and intense pain conditions [9]. In contrast, the outliers of DF(LASSO) do not exhibit an easily recognizable pattern (Figure 3(B)), consisting of both low and high pain scores. This may be due to its limited ability to detect multiple outliers, causing challenges for meaningful interpretation and reliable inference.
6.2 Effects of Influential Observations on Variable Selection
The detected influential points demonstrate non-negligible impact on variable selection. For ClusMIP, this is assessed by computing the average size of submodels selected from the following three datasets: the full dataset and the two reduced datasets upon removing $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(Selector)}}$ and $\widehat{\text{I}}^{\text{infl}}_{\text{Common}}$; see Figure 4(A). For DF(LASSO), equivalent results, shown in Figure 5(A), are obtained from the full dataset and only one reduced dataset obtained by removing $\widehat{\text{I}}^{\text{infl}}_{\text{DF(LASSO)}}$ since $\widehat{\text{I}}^{\text{infl}}_{\text{Common}}$ does not overlap with $\widehat{\text{I}}^{\text{infl}}_{\text{DF(LASSO)}}$.
In Panel (A) of Figures 4 and 5, variation in the average model sizes before and after outlier removal is clearly seen. Specifically, such quantities for LASSO, ENET, SCAD and MCP before outlier removal via ClusMIP are 83.63, 367.72, 32.95 and 18.87, which change to 99.61, 379.78, 21.51 and 13.16 after the removal. Moreover, for LASSO, the model size increases from 83.63 to 91.65 upon removing those identified via DF(LASSO). Furthermore, while the model sizes increase after outlier removal for the two convex penalties LASSO and ENET, those selected by the two non-convex penalties SCAD and MCP are more sparse after removal of detected outliers. Also, LASSO and ENET select considerably more predictors than SCAD and MCP both before and after eliminating identified influential observations.
To supplement Panels (A) of Figures 4 and 5, we also calculate the average number of common predictors that are contained in submodels selected from both full and reduced datasets, and those that are unique to the submodels chosen from the full and reduced datasets, shown in Panels (B) of Figures 4 and 5, respectively. From these plots, a significant number of predictors unique only to the reduced datasets are seen for all model selectors. Specifically, 58.99, 84.04, 16.22 and 12.00 unique predictors are chosen by LASSO, ENET, SCAD and MCP based on the reduced datasets obtained by removing $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(Selector)}}$, consisting of $59.22\%$, $22.12\%$, $75.40\%$ and $91.19\%$ of their respective submodels. Moreover, LASSO selects 39.60 unique predictors upon removing $\widehat{\text{I}}^{\text{infl}}_{\text{DF(LASSO)}}$, $43.21\%$ of its submodel size. These phenomena further indicates the impact of influential observations on variable selection.
@@ 6.2 Effects of Influential Observations on Variable Selection
In Figure 6, we illustrate the effects of influential points on variable selection by highlighting the brain regions selected via MCP with color intensities represented by the corresponding coefficient estimates obtained from the full and reduced datasets. Panels A and B show the discrepancy in selected regions, while Panels C, D and E pinpoint common and uniquely selected regions from the two datasets. Six regions are common to both models, including primary visual, frontoparietal, dorsal attention, and default mode areas, as well as the amygdala. Regions unique to the full model include the cerebellum, while regions unique to the reduced model include ventral attention and additional frontoparietal areas. The frontoparietal network, dorsal attention network, visual networks and the amygdala are all associated with pain processing, justifying their inclusion in both models. On the other hand, the ventral attention network (VAN) is responsible for processing unexpected external stimuli. In this study, stronger responses resulting from the high pain scores may lead to more robust bottom-up activation of the VAN. This justifies the inclusion of the VAN in the model chosen from the reduced dataset after removal of the “outlying” low pain scores.
Figure 4: ClusMIP. (A): average sizes of submodels selected from the full and reduced datasets upon removing $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(Selector)}}$ and $\widehat{\text{I}}^{\text{infl}}_{\text{Common}}$. (B): average number of predictors included in submodels selected from both full and reduced datasets, unique to those from the full dataset, and unique to those from the reduced dataset. (C): change in the normed coefficient estimates upon outlier removal. (D): sample correlation between observed and predicted pain scores. (E): $-\log{(\text{MSE of Prediction})}$.
Figure 5: DF(LASSO). (A): average size of submodels chosen from the full and reduced datasets (removing $\widehat{\text{I}}^{\text{infl}}_{\text{DF(LASSO)}}$). (B): average number of predictors that are in the submodels selected from both full and reduced datasets, unique to those from the full dataset, and unique to those from the reduced dataset. (C): sample correlation between observed and predicted pain scores. (D): $-\log{(\text{MSE of Prediction})}$.
6.3 Effects of Influential Observations on Coefficient Estimation
In Figure 4(C), we examine the impact of influential points on coefficient estimation by computing the percentage change in the $\ell_{2}$-normed coefficient estimates defined by $(||\widehat{\boldsymbol{\beta}}_{\text{full}}||_{2}-||\widehat{\boldsymbol{\beta}}_{\text{redu}}||_{2})/||\widehat{\boldsymbol{\beta}}_{\text{full}}||_{2}$, where $\widehat{\boldsymbol{\beta}}_{\text{full}}$ and $\widehat{\boldsymbol{\beta}}_{\text{redu}}$ are the coefficient estimates obtained by fitting the full dataset and reduced dataset via removing $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(Selector)}}$ or $\widehat{\text{I}}^{\text{infl}}_{\text{Common}}$. From Figure 4(C), at least $30\%$ change in the $\ell_{2}$-normed coefficients is seen for all model selectors from these two reduced datasets. The largest changes, $63.03\%$ and $82.94\%$, are observed for SCAD. In comparison, although LASSO shows the smallest coefficient adjustments, they reach significant levels of $33.11\%$ and $31.53\%$. The same benchmark is also calculated for DF(LASSO) upon removing $\widehat{\text{I}}^{\text{infl}}_{\text{DF(LASSO)}}$ only, which attains $34.22\%$. These results highlight the considerable impact of outliers on coefficient estimation.
6.4 Effects of Influential Observations on Prediction
Improved prediction power is also observed upon removal of detected influential observations. This is assessed using two criteria: the sample correlation between observed and predicted pain scores and the logarithm of MSE for prediction. They are shown separately in Figure 4(D) and (E) for ClusMIP, and Figure 5(C) and (D) for DF(LASSO). Clear improvements are seen after removing detected influential observations for all model selectors. For example, the highest sample correlation is obtained by MCP, which increases from 0.53 to 0.80 after removing $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(MCP)}}$. This is accompanied by improved performance in the MSE of prediction, which declines from $\exp{(9.32)}=11203.49$ to $\exp{(8.79)}=6542.97$, the largest such reduction ($41.6\%$) among all model selectors. In contrast, no significant difference of sample correlation is seen for LASSO upon removing the outliers detected by DF(LASSO) (Figure 5(C)). In contrast, the MSE of LASSO drops from $\exp{(9.16)}=9551.72$ to $\exp{(9.10)}=8916.39$ after removing the outliers detected by DF(LASSO), which is $36.27\%$ higher than that obtained after removing the outliers detected by ClusMIP(MCP). Thus, the difference in prediction power before and after outlier elimination illustrates the benefits of carefully considering points identified by all detection procedures. Moreover, the advantage of ClusMIP over DF(LASSO) in terms of predictive performance is clearly observed.
@@ 6.4 Effects of Influential Observations on Prediction
6.5 Summary of Analysis
In developing a neurologic signature of physical pain, the principal objective is to create a single model predictive of pain levels. It is critical to remove influential observations specific to the particular model selector used for deriving the model. In the context of our thermal pain data, the influential observations detected by ClusMIP(MCP) correspond to low pain scores (Figure 3(A)) for each temperature. This pattern highlights the distinction between low and intense pain processing mechanisms, consistent with findings in other pain studies. Moreover, after exclusion of these points, MCP selects a statistical model containing unique feature(s) (the VAN network) (Figure 6), which regulates the responses corresponding to high pain scores. Furthermore, the reduced model selected by MCP is the most sparse (Figure 4(A)), yet delivers the best prediction power (Figure 4(D) and (E)) among all competing selectors. Specifically, it demonstrates the highest correlation between observed and predicted pain scores, and the lowest MSE in prediction. In other words, upon removing $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(MCP)}}$, MCP selects the most parsimonious, scientifically justifiable, and interpretable model while exhibiting the highest predictive performance. Therefore, we posit that the best performing detection procedure for the thermal pain dataset is ClusMIP(MCP).
7 Conclusion and Further Discussion
In this manuscript, we have described the existence of multiple factors that can negatively influence the development of a neurologic signature of physical pain [21], including inter-subject variability in pain processing and functional topology, and the noisy nature of fMRI data. Together these factors can hinder the formation of a generalizable predictive model, leading to concerns about the validity of subsequent statistical inference. We illustrate how these factors can unduly affect the choice of a stochastically selected submodel, an issue that is largely ignored in the existing influential diagnosis literature. In our simulation study, we observe that failure to account for outliers leads to non-negligible, detrimental impacts on various aspects of downstream statistical analysis, including distorted variable selection outcomes, biased coefficient estimates and worse prediction performance. One existing method DF(LASSO) falls short in its theoretical foundation and is not designed to identify multiple influential observations. The inability to detect a block of outliers also holds for the other existing method MIP under certain common contamination scenarios. As such, the existence of outliers in fMRI data, the pitfalls of ignoring them, and the limitations of existing methods necessitate the development of a theoretically sound detection method capable of identifying multiple influential points.
To guide the development of our detection method, we first recognize that variable selection is pervasive across an extensive scope of scientific disciplines for both explanatory and predictive modeling. Moreover, different high-dimensional model selectors (i.e., penalties) have distinctive theoretical properties designed to handle different data types. In fact, in fMRI research, regularized model selectors with both convex and non-convex penalties have been widely applied. Thus, although the motivation for this work originates from the aforementioned thermal pain study, we target a universally adaptable detection framework applicable well beyond the specific application. Consequently, we introduce the following set of guiding principles: (i) the ability to detect multiple influential observations; (ii) applicability to different model selectors and data types; and (iii) theoretical justification.
One natural question based on our guiding principles is how to interpret identified influential points, as different model selectors may well lead to different sets of outliers. We believe this ultimately depends on the situation. Therefore, we deliberately put forward a spectrum of detection approaches corresponding to different model selectors. The choice of one or multiple model selector(s) hinges on elements which include, but are not limited to, the purpose of study and domain science expertise. Consequently, it is the practitioner’ prerogative to decide on the final influential points, ranging from the intersection to the union of all the possible influential point sets. In our case, the goal is to determine a model predictive of pain scores. Precise instruction on the model selector and detection procedure is offered based on the interpretability of detected data points and reduced model, and its resulting performance.
Using these guiding principles, the GDF measure was proposed in (3), and its exchangeability and CLT established in Theorems 1 and 2. Combination of the GDF and consistent high-dimensional clustering technique leads to our ClusMIP procedure (Algorithm ClusMIP), which controls the FPR in both finite- and large-sample (Theorem 3). Based on the simulations in Section 5, the ClusMIP procedure demonstrates a clear advantage over the existing methods (DF(LASSO) and MIP) in terms of improved detection power, leading to considerably enhanced prediction power (lower MSE in prediction) and variable selection performance (higher empirical probability in selecting a submodel containing the true model) after removal of detected outliers. Specifically, the detection power of ClusMIP(SLASSO) is consistently performs best, varying between 0.90 and 1.00. In contrast, under-performance is seen for DF(LASSO) and MIP, where the detection power reaches levels between 0 and 0.20.
Applying the methods to the pain data, we find that the identified outliers from the ClusMIP procedure uniformly exhibit a tractable pattern corresponding to low pain ratings (Figure 3(A)). This substantiates the discrepancy between low and intense pain processing conditions. In contrast, those identified by DF(LASSO) show no recognizable pattern. The reduced model, obtained by removing observations detected by ClusMIP(MCP), is scientifically interpretable, containing unique features (i.e., the VAN network) that regulates the neuronal responses from the remaining high pain scores (Figure 6). In addition, the impact of excluding detected outliers on variable selection and coefficient estimation is clearly shown for all model selectors and detection procedures. Specifically, notable variation in the average model size and normed coefficient estimates (at least $30\%$ change) are seen (Figure 4(A) to (C), Figure 5(A) and (B)). Remarkably, eliminating the outliers detected by ClusMIP(MCP) leads to the highest prediction power attained among all competing methods (Figure 4(D) and (E)), with corresponding MSE $26.62\%$ lower than those obtained using DF(LASSO) (Figure 5(C) and (D)). Since MCP selects sparsest submodel with plausible scientific interpretation and highest prediction performance after removal of outliers identified by ClusMIP(MCP), it is recommended among all the competing proposals for this application.